### v2

@@ 5.3 Strategies and play decompositions
We first introduce the general idea of a strategy based on a finite simple decomposition of a play in Section 5.3.1. We then partially define Mealy machines that induce strategies based on finite simple decompositions in Section 5.3.2.
@@ 5.3.1 Strategies based on a play decomposition
First, we define the histories for which players continue following the current segment of the play. We call such histories coherent with $\mathcal{D}$. Coherence is defined inductively: we say that a history is coherent with $\mathcal{D}$ if there is some $j\in\llbracket{}k\rrbracket{}$ such that it is $j$-coherent with $\mathcal{D}$. The base case of the induction is the history $v_{0}$; it is $1$-coherent with $\mathcal{D}$. We now consider a $j$-coherent history $h$ and let $v\in\mathsf{Succ}_{E}(\mathsf{last}(h))$. If $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$, then $hv$ is $(j+1)$-coherent. Otherwise, if $v$ occurs in $\mathsf{sg}_{j}$, then $hv$ is $j$-coherent. In any other case, $hv$ is not coherent with $\mathcal{D}$.
We now define strategies based on $\mathcal{D}$, i.e., strategies that, given a coherent history, attempt to complete the segment in progress. Given a history $h$ that is $j$-coherent, we call the vertex that follows $\mathsf{last}(h)$ in $\mathsf{sg}_{j}$ the next vertex of $h$ (with respect to $\mathcal{D}$). We prove that this vertex is well-defined.
@@
Let $h$ be a history that is coherent with the simple decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k})$. The next vertex of $h$ (with respect to $\mathcal{D}$) is well-defined.
Proof 5.7.
@@
We therefore assume that $j<k$ and show the existence of a next vertex by contradiction. Assume there is no suitable vertex. It implies that $\mathsf{last}(h)=\mathsf{last}(\mathsf{sg}_{j})$. By simplicity and the absence of trivial histories in a decomposition, we have $\mathsf{first}(h_{j})\neq\mathsf{last}(h_{j})$. Therefore, there must be a prefix of $h$ that is $j$-coherent by definition of $j$-coherence. We obtain that $h$ should either be $(j+1)$-coherent or not coherent, a contradiction.
We say that a strategy of $\mathcal{P}_{i}$ is based on $\mathcal{D}$ if to any history $h\in\mathsf{Hist}_{i}(\mathcal{A})$ that is coherent with $\mathcal{D}$, it assigns the next vertex of $h$ with respect to $\mathcal{D}$. Any strategy profile $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ such that $\sigma_{i}$ is based on $\mathcal{D}$ for all $i\in\llbracket{}n\rrbracket{}$ is such that $\mathsf{Out}(\sigma,v_{0})=\pi$.
@@ 5.3.2 Finite-memory decomposition-based strategies
Let $\pi=v_{0}v_{1}\ldots\in\mathsf{Plays}(\mathcal{A})$ be a play that admits a simple segment decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k})$. We partially define a Mealy machine that serves as the basis for the finite-memory NEs described in the next section.
The memory state space is made of pairs of the form $(\mathcal{P}_{i},j)$ for some $j\in\llbracket{}k\rrbracket{}$. We do not consider all such pairs, e.g., it is not necessary in Example 5.1. Therefore, we parameterise our construction by a non-empty set of players $I\subseteq\llbracket{}n\rrbracket{}$. We consider the memory state space $M^{I,\mathcal{D}}=\{\mathcal{P}_{i}\mid i\in I\}\times\llbracket{}k\rrbracket{}$. The initial state $m^{I,\mathcal{D}}_{\mathsf{init}}$ is any state of the form $(\mathcal{P}_{i},1)\in M^{I,\mathcal{D}}$.
The update function $\mathsf{up}^{I,\mathcal{D}}$ behaves similarly to Figure 3(b). It keeps track of the last player in $I$ to have moved and the current segment. Formally, for any $(\mathcal{P}_{i},j)\in M^{I,\mathcal{D}}$ and vertex $v$ occurring in $\mathsf{sg}_{j}$, we let $\mathsf{up}^{I,\mathcal{D}}((\mathcal{P}_{i},j),v)=(\mathcal{P}_{i^{\prime}},j^{\prime})$ where (i) $i^{\prime}$ is such that $v\in V_{i^{\prime}}$ if $v\in\bigcup_{i^{\prime\prime}\in I}V_{i^{\prime\prime}}$ and otherwise $i^{\prime}=i$, and (ii) $j^{\prime}=j+1$ if $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$ and $j^{\prime}=j$ otherwise. Updates from $(\mathcal{P}_{i},j)$ for a vertex that does not appear in $\mathsf{sg}_{j}$ are left undefined.
The next-move function $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ of $\mathcal{P}_{i}$ proposes the next vertex of the current segment. Formally, given a memory state $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and a vertex $v\in V_{i}$ that occurs in $\mathsf{sg}_{j}$, we let $\mathsf{nxt}_{i}^{I,\mathcal{D}}((\mathcal{P}_{i^{\prime}},j),v)$ be the vertex occurring after $v$ in $\mathsf{sg}_{j+1}$ if $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$, and otherwise we let it be the vertex occurring after $v$ in $\mathsf{sg}_{j}$. Like updates, the next-move function is left undefined in memory states $(\mathcal{P}_{i},j)$ for a vertex that does not appear in $\mathsf{sg}_{j}$.
@@
Let $\pi$ be a play that admits a simple segment decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k})$. Let $I\subseteq\llbracket{}n\rrbracket{}$ be a non-empty set. We prove that any finite-memory strategy induced by a Mealy machine that extends the partially defined Mealy machine $(M^{I,\mathcal{D}},m^{I,\mathcal{D}}_{\mathsf{init}},\mathsf{up}^{I,\mathcal{D}},\mathsf{nxt}_{i}^{I,\mathcal{D}})$ is based on $\mathcal{D}$. To this end, we establish that if a history $h$ is $j$-coherent with $\mathcal{D}$, then the memory state after the Mealy machine reads $h$ is of the form $(\mathcal{P}_{i},j)$.
@@
Let $i\in\llbracket{}n\rrbracket{}$. Let $\mathcal{M}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ be a Mealy machine of $\mathcal{P}_{i}$ such that $M^{I,\mathcal{D}}\subseteq M$, $m^{I,\mathcal{D}}_{\mathsf{init}}=m_{\mathsf{init}}$, $\mathsf{up}$ and $\mathsf{nxt}$ coincide with $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ respectively on the domain of the latter functions. The strategy $\sigma_{i}$ induced by $\mathcal{M}$ is based on $\mathcal{D}$ and for all $h\in\mathsf{Hist}(\mathcal{A})$, if $h$ is $j$-coherent with $\mathcal{D}$, then $\widehat{\mathsf{up}}(h)=(\mathcal{P}_{i^{\prime}},j)$ for some $i^{\prime}\in I$.
Proof 5.9.
We first show the second claim of the lemma. We proceed by induction on the number of vertices in $h$. The only coherent history with a single vertex is $v_{0}$. Our assumption on $\mathcal{D}$ ensures that $v_{0}\neq\mathsf{last}(\mathsf{sg}_{1})$, thus we have that $\mathsf{up}(m_{\mathsf{init}},v_{0})$ is of the form $(\mathcal{P}_{i^{\prime}},1)$.
@@ 5.4 Nash equilibria in reachability games
We show that from any NE in a reachability game, we can derive a finite-memory NE from the same initial vertex in which the same targets are visited. We build on the NE outcomes with simple decompositions provided by Lemma 5.3 and extend the Mealy machines of Section 5.3.2 to obtain NEs with a memory size of at most $n^{2}$. We fix a reachability game $\mathcal{G}=(\mathcal{A},(\mathsf{Reach}(T_{i}))_{i\in\llbracket{}n\rrbracket{}})$.
The general idea of the construction is to use the memory state space $M^{I,\mathcal{D}}$ where $I\subseteq\llbracket{}n\rrbracket{}$ is the set of players who do not see their targets if it is non-empty, or a single arbitrary player if all players see their target. Let $i^{\prime}\in I$ and $j\in\llbracket{}k\rrbracket{}$. We extend $\mathsf{up}^{I,\mathcal{D}}$ to leave unchanged the memory state if in state $(\mathcal{P}_{i^{\prime}},j)$ whenever the current vertex is not in $\mathsf{sg}_{j}$. In this same situation, for $i\neq i^{\prime}$, the next-move function $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ is extended to assign moves from a memoryless uniformly winning strategy of the second player in the coalition game $\mathcal{G}_{i^{\prime}}=(\mathcal{A}_{i^{\prime}},\mathsf{Reach}(T_{i^{\prime}}))$ (which exists by Theorem 3.1), and is left arbitrary if $i=i^{\prime}$. The equilibrium’s stability is a consequence of Theorem 4.1 and Lemma 3.2: if the target of $\mathcal{P}_{i}$ is not visited in the intended outcome, all vertices along this play are not winning for the first player of the coalition game $\mathcal{G}_{i}$. We formalise the explanation above in the proof of the following theorem.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$ in $\mathcal{G}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))=\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ where each strategy of $\sigma$ has a memory size of at most $n^{2}$. More precisely, we can bound their memory size by
$\max\{1,n-|\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))|\}\cdot\max\{1,|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|\}.$
Proof 5.11.
@@
We formally extend the Mealy machines of Section 5.3.2. Let $i\in\llbracket{}n\rrbracket{}$. We consider the Mealy machine $\mathcal{M}_{i}=(M^{I,\mathcal{D}},m^{I,\mathcal{D}}_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ where $\mathsf{up}$ and $\mathsf{nxt}_{i}$ respectively extend $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ as follows. Let $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$. We let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=(\mathcal{P}_{i^{\prime}},j)$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and otherwise we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be arbitrary. We let $\sigma_{i}$ denote the strategy induced by $\mathcal{M}_{i}$. It is easy to see that the memory size of $\sigma_{i}$ is at most $n^{2}$. It follows from Lemma 5.8 that the outcome of $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ from $v_{0}$ is $\pi$.
We now prove that $\sigma$ is an NE from $v_{0}$. It suffices to show that for all $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$, $\mathcal{P}_{i}$ does not have a profitable deviation. Fix $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$. We show that all histories that are consistent with the strategy profile $\sigma_{-i}=(\sigma_{i^{\prime}})_{i^{\prime}\neq i}$ do not leave $W_{-i}(\mathsf{Safe}(T_{i}))$. We proceed by induction on the length of histories that start in $v_{0}$ and are consistent with $\sigma_{-i}$. To establish the property above, we show in parallel that for any such history $hv$, if $(\mathcal{P}_{i^{\prime}},j)=\widehat{\mathsf{up}}(h)$, we have the implication: if $v$ does not occur in $\mathsf{sg}_{j}$, then $i^{\prime}=i$.
@@
If there exists an NE from $v_{0}$ such that no (resp. all) players see their target in its outcome, then there is a finite-memory NE from $v_{0}$ such that no (resp. all) players see their target in its outcome such that all strategies have a memory size of at most $n$.

### v3

@@ 5.3 Strategies and play decompositions
We first introduce the general idea of a strategy based on a finite simple decomposition of a play in Section 5.3.1. We then define Mealy machines that induce strategies based on finite simple decompositions in Section 5.3.2.
@@ 5.3.1 Strategies based on a play decomposition
First, we define the histories for which players continue following the current segment of the play. We call such histories coherent with $\mathcal{D}$.
Definition 5.6 (History coherent with a decomposition).
A history is coherent with $\mathcal{D}$ if there is some $j\in\llbracket{}k\rrbracket{}$ such that it is $j$-coherent with $\mathcal{D}$. We define $j$-coherence inductively. The only one-vertex history that is $1$-coherent with $\mathcal{D}$ is the history $h=v_{0}$. For all $h\in\mathsf{Hist}(\mathcal{A})$ and $v\in\mathsf{Succ}_{E}(\mathsf{last}(h))$, we define that
  • if $h$ is $j$-coherent for some $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$, then $hv$ is $(j+1)$-coherent;
  • otherwise, if $h$ is $j$-coherent for some $j\in\llbracket{}k\rrbracket{}$ and $v$ occurs in $\mathsf{sg}_{j}$, then $hv$ is $j$-coherent;
  • in any other case, $hv$ is not coherent with $\mathcal{D}$.
As the cases in the above definition are disjoint, given a history $h$ coherent with $\mathcal{D}$, there is a unique $j\in\llbracket{}k\rrbracket{}$ such that $h$ is $j$-coherent with $\mathcal{D}$.
We now define strategies based on $\mathcal{D}$, i.e., strategies that, given a coherent history, attempt to complete the segment in progress. We use the following notion in this definition.
Definition 5.7.
Let $h$ be a history that is $j$-coherent with $\mathcal{D}$. The $\mathcal{D}$-next vertex of $h$ to be the vertex that follows $\mathsf{last}(h)$ in $\mathsf{sg}_{j}$.
Given a $j$-coherent history $h$, its $\mathcal{D}$-next vertex is well-defined only if $\mathsf{last}(h)$ differs from $\mathsf{last}(\mathsf{sg}_{j})$. (note that uniqueness follows from $\mathcal{D}$ being simple). We show that this is always the case below.
@@
Let $h$ be a history that is coherent with the simple decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k})$. The $\mathcal{D}$-next vertex of $h$ is well-defined.
Proof 5.9.
@@
We therefore assume that $j<k$ and show the existence of a $\mathcal{D}$-next vertex by contradiction. Assume there is no suitable vertex. It implies that $\mathsf{last}(h)=\mathsf{last}(\mathsf{sg}_{j})$. By simplicity and the absence of trivial histories in $\mathcal{D}$, we have $\mathsf{first}(\mathsf{sg}_{j})\neq\mathsf{last}(\mathsf{sg}_{j})$. Therefore, there must be a prefix of $h$ that is $j$-coherent by definition of $j$-coherence. We obtain that $h$ should either be $(j+1)$-coherent or not coherent, a contradiction.
We can now formalise the notion of a strategy based on $\mathcal{D}$.
Definition 5.10.
A strategy of $\mathcal{P}_{i}$ is based on $\mathcal{D}$ if to any history $h\in\mathsf{Hist}_{i}(\mathcal{A})$ that is coherent with $\mathcal{D}$, it assigns the $\mathcal{D}$-next vertex of $h$.
We now formally show below that if all players use strategies that are based on $\mathcal{D}$ from $\mathsf{first}(\pi)$, then the resulting outcome is $\pi$.
Lemma 5.11.
Let $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ be a strategy profile such that $\sigma_{i}$ is based on $\mathcal{D}$ for all $i\in\llbracket{}n\rrbracket{}$. Then $\mathsf{Out}(\sigma,\mathsf{first}(\pi))=\pi$.
Proof 5.12.
We consider two cases. First, assume that $\mathcal{D}=(\pi)$. It follows that all prefixes of $\pi$ are $1$-coherent with $\mathcal{D}$, and the claim follows from the definition of $\mathcal{D}$-next vertex.
Assume now that $k>1$. We let $\pi=v_{0}v_{1}\ldots$ and $\ell_{1}<\ldots<\ell_{k-1}$ be indices such that, for all $j\in\llbracket{}k-1\rrbracket{}$, $\pi{\ell_{j}}=\mathsf{sg}_{1}\cdot\ldots\cdot\mathsf{sg}_{j}$. To end the proof, due to the definition of a $\mathcal{D}$-next vertex, it suffices to show that for all $\ell\in\mathbb{N}$:
  • if $\ell<\ell_{1}$, then $\pi{\ell}$ is $1$-coherent with $\mathcal{D}$,
  • if $\ell_{j-1}\leq\ell<\ell_{j}$ for some $1<j<k$, then $\pi{\ell}$ is $j$-coherent with $\mathcal{D}$,
  • if $\ell\geq\ell_{k-1}$, then $\pi{\ell}$ is $k$-coherent with $\mathcal{D}$.
We prove the above points by induction. For $\ell=0$, the $1$-coherence of $\pi{0}=v_{0}$ is due to the base case of the definition of coherence. Now assume by induction that the claim holds for $\ell\in\mathbb{N}$ and let us show it for $\ell+1$.
First, let us assume that $\ell+1<\ell_{1}$. The $1$-coherence of $\pi{\ell+1}$ follows from the $1$-coherence of $\pi{\ell}$ and the fact that $v_{\ell+1}$ occurs in $\mathsf{sg}_{1}$ and that $v_{\ell+1}\neq\mathsf{last}(\mathsf{sg}_{1})=v_{\ell_{1}}$ by simplicity of $\mathsf{sg}_{1}$.
We now assume that there exists some $1<j<k$ such that $\ell_{j-1}\leq\ell+1<\ell_{j}$. We distinguish two cases. If $\ell+1=\ell_{j-1}$, then by induction, $\pi{\ell}$ is (j1)(j-1)-coherent and the $j$-coherence of $\pi{\ell}$ follows from $v_{\ell+1}=\mathsf{last}(\mathsf{sg}_{j-1})$. If $\ell+1>\ell_{j-1}$, then by induction, $\pi{\ell}$ is $j$-coherent; the $j$-coherence of $\pi{\ell}$ is implied by $v_{\ell+1}$ occurring in $\mathsf{sg}_{j}$ and $v_{\ell+1}\neq\mathsf{last}(\mathsf{sg}_{j})=v_{\ell_{j}}$ (which follows from $\mathcal{D}$ being simple).
The case $\ell\geq\ell_{k-1}$ can be handled in the same way as the previous one, and is omitted to limit repetition.
@@ 5.3.2 Finite-memory decomposition-based strategies
We generalise the common components of the Mealy machines used in Examples 5.1 and 5.2. For the remainder of the section, we let $\pi=v_{0}v_{1}\ldots\in\mathsf{Plays}(\mathcal{A})$ be a play that admits a finite simple segment decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k})$ that has no trivial segments.
We consider Mealy machines that, along coherent histories, keep track of two pieces of information: a player and a segment index. This is achieved using memory states of the form $(\mathcal{P}_{i},j)$. We do not require all such pairs, e.g., it is not necessary in Example 5.1, and instead we parameterise our definition by a non-empty set of players $I\subseteq\llbracket{}n\rrbracket{}$. Our construction is based on Mealy machines whose memory state space includes $M^{I,\mathcal{D}}=\{\mathcal{P}_{i}\mid i\in I\}\times\llbracket{}k\rrbracket{}$. There may be additional memory states, as is the case in Example 5.2. We require that the initial state of the Mealy machine be of the form $(\mathcal{P}_{i},1)$ for some $i\in I$.
Intuitively, the memory of these Mealy machines keeps track of the index of the current segment of the play (incrementing it when the last vertex of the current segment is reached) and the last player in $I$ to have moved. We define the next-move function so that it proposes the $\mathcal{D}$-next vertex after a history coherent with $\mathcal{D}$. If some vertex outside of the current segment is processed (which only happens when dealing with histories that are not coherent with $\mathcal{D}$), we do not impose any constraint on the behaviour of the update and next-move functions. We generalise this aspect differently depending on whether we consider reachability games or shortest-path games in the sequel.
We formalise the information explanation above as follows.
Definition 5.13 ($(I,\mathcal{D})$-compatible Mealy machines).
Let $I\subseteq\llbracket{}n\rrbracket{}$. We define $M^{I,\mathcal{D}}=\{\mathcal{P}_{i}\mid i\in I\}\times\llbracket{}k\rrbracket{}$. We let $\mathsf{up}^{I,\mathcal{D}}\colon M^{I,\mathcal{D}}\times V\to M$ denote the (partial) function such that for all $(\mathcal{P}_{i},j)\in M^{I,\mathcal{D}}$ and all vertices $v$ occurring in $\mathsf{sg}_{j}$, $\mathsf{up}^{I,\mathcal{D}}((\mathcal{P}_{i},j),v)=(\mathcal{P}_{i^{\prime}},j^{\prime})$ such that
  1. (a)
    $i^{\prime}$ is such that $v\in V_{i^{\prime}}$ if $v\in\bigcup_{i^{\prime\prime}\in I}V_{i^{\prime\prime}}$ and otherwise $i^{\prime}=i$,
  2. (b)
    $j^{\prime}=j+1$ if $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$ and $j^{\prime}=j$ otherwise.
For a given $i\in\llbracket{}n\rrbracket{}$, we let $\mathsf{nxt}_{i}^{I,\mathcal{D}}\colon M^{I,\mathcal{D}}\times V\to V$ denote the (partial) function such that for all $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and vertices $v\in V_{i}$ that occur in $\mathsf{sg}_{j}$, $\mathsf{nxt}_{i}^{I,\mathcal{D}}((\mathcal{P}_{i^{\prime}},j),v)=v^{\prime}$ such that
  1. (a)
    if $j<k$ and $v=\mathsf{last}(\mathsf{sg}_{j})$, then $v^{\prime}$ is the vertex occurring after $v$ in $\mathsf{sg}_{j+1}$,
  2. (b)
    otherwise, we let $v^{\prime}$ be the vertex occurring after $v$ in $\mathsf{sg}_{j}$.
A Mealy machine $\mathcal{M}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ of $\mathcal{P}_{i}$ is $(I,\mathcal{D})$-compatible if $M^{I,\mathcal{D}}\subseteq M$, $m_{\mathsf{init}}$ is of the form $(\mathcal{P}_{i},1)$ for some $i\in I$, and $\mathsf{up}$ and $\mathsf{nxt}_{i}$ coincide with $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ respectively on the domain of the latter functions.
@@
Let $I\subseteq\llbracket{}n\rrbracket{}$ be a non-empty set. We prove that any finite-memory strategy induced by an $(I,\mathcal{D})$-compatible Mealy machine is based on $\mathcal{D}$. To this end, we establish that if a history $h$ is $j$-coherent with $\mathcal{D}$, then the memory state after the Mealy machine reads $h$ is of the form $(\mathcal{P}_{i},j)$.
@@
Let $i\in\llbracket{}n\rrbracket{}$. Let $\mathcal{M}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ be a Mealy machine of $\mathcal{P}_{i}$ that is $(I,\mathcal{D})$-compatible. The strategy $\sigma_{i}$ induced by $\mathcal{M}$ is based on $\mathcal{D}$ and for all $h\in\mathsf{Hist}(\mathcal{A})$, if $h$ is $j$-coherent with $\mathcal{D}$, then $\widehat{\mathsf{up}}(h)=(\mathcal{P}_{i^{\prime}},j)$ for some $i^{\prime}\in I$.
Proof 5.15.
We first show the second claim of the lemma. We proceed by induction on the number of vertices in $h$. The only coherent history with a single vertex is $v_{0}$. If $\mathsf{sg}_{1}=\pi$ (i.e., $\mathcal{D}$ is a trivial decomposition), then $\mathsf{up}(m_{\mathsf{init}},v_{0})$ is of the form $(\mathcal{P}_{i^{\prime}},1)$ (the definition of $\mathsf{up}^{I,\mathcal{D}}$ in Definition 5.13 has only one case). Assume now that $\mathcal{D}$ is a non-trivial decomposition. It follows that $\mathsf{sg}_{1}$ is a non-trivial simple history, and thus $v_{0}\neq\mathsf{last}(\mathsf{sg}_{1})$, By definition of $\mathsf{up}^{I,\mathcal{D}}$, $\mathsf{up}(m_{\mathsf{init}},v_{0})$ is again of the form $(\mathcal{P}_{i^{\prime}},1)$.
@@ 5.4 Nash equilibria in reachability games
We show that from any NE in a reachability game, we can derive a finite-memory NE from the same initial vertex in which the same targets are visited. We build on the NE outcomes with simple decompositions provided by Lemma 5.3 and the Mealy machines of Definition 5.13 to obtain NEs with a memory size of at most $n^{2}$. We fix a reachability game $\mathcal{G}=(\mathcal{A},(\mathsf{Reach}(T_{i}))_{i\in\llbracket{}n\rrbracket{}})$.
Let $I\subseteq\llbracket{}n\rrbracket{}$ is the set of players who do not see their targets if it is non-empty, or a single arbitrary player if all players see their target. The general idea of the construction is to use, for each $i\in\llbracket{}n\rrbracket{}$, an $(I,\mathcal{D})$-compatible Mealy machine $\mathcal{M}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ with the memory state space $M=M^{I,\mathcal{D}}$. Let $i^{\prime}\in I$ and $j\in\llbracket{}k\rrbracket{}$. We define $\mathsf{up}$ such that it leaves the memory state unchanged from a state $(\mathcal{P}_{i^{\prime}},j)$ whenever an update is made with a vertex that does not occur in $\mathsf{sg}_{j}$. In this same situation, if $i\neq i^{\prime}$, $\mathsf{nxt}_{i}$ assign moves from a memoryless uniformly winning strategy of the second player in the coalition game $\mathcal{G}_{i^{\prime}}=(\mathcal{A}_{i^{\prime}},\mathsf{Reach}(T_{i^{\prime}}))$ (which exists by Theorem 3.1), and is left arbitrary if $i=i^{\prime}$. The equilibrium’s stability is a consequence of Theorem 4.1 and Lemma 3.2: if the target of $\mathcal{P}_{i}$ is not visited in the intended outcome, all vertices along this play are not winning for the first player of the coalition game $\mathcal{G}_{i}$. We formalise the explanation above in the proof of the following theorem.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$ in $\mathcal{G}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))=\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ where each strategy of $\sigma$ has a memory size of at most
$\max\{1,n-|\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))|\}\cdot\max\{1,|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|\}\leq n^{2}.$
Proof 5.17.
@@
Let $i\in\llbracket{}n\rrbracket{}$. We define an $(I,\mathcal{D})$-compatible Mealy machine $\mathcal{M}_{i}=(M^{I,\mathcal{D}},(\mathcal{P}_{\min I},1),\mathsf{up},\mathsf{nxt}_{i})$ for $\mathcal{P}_{i}$ as follows (cf. Definition 5.13). The functions $\mathsf{up}$ and $\mathsf{nxt}_{i}$ respectively extend $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ as follows. Let $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$. We let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=(\mathcal{P}_{i^{\prime}},j)$. If $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and, let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be arbitrary, otherwise.
We let $\sigma_{i}$ denote the strategy induced by $\mathcal{M}_{i}$. It is easy to see that the memory size of $\sigma_{i}$ is at most $n^{2}$. It follows from Lemma 5.14 that the outcome of $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ from $v_{0}$ is $\pi$.
We now prove that $\sigma$ is an NE from $v_{0}$. It suffices to show that for all $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$, $\mathcal{P}_{i}$ does not have a profitable deviation. Fix $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$. We show that all histories that are consistent with the strategy profile $\sigma_{-i}=(\sigma_{i^{\prime}})_{i^{\prime}\neq i}$ do not leave $W_{-i}(\mathsf{Safe}(T_{i}))$. We proceed by induction on the length of histories that start in $v_{0}$ and are consistent with $\sigma_{-i}$. To establish the property above, we show in parallel that for any such history $hv$, letting $(\mathcal{P}_{i^{\prime}},j)=\widehat{\mathsf{up}}(h)$, we have the implication: if $v$ does not occur in $\mathsf{sg}_{j}$, then $i^{\prime}=i$.
@@
If there exists an NE from a vertex $v_{0}$ such that no (resp. all) players see their target in its outcome, then there is a finite-memory NE from $v_{0}$ such that no (resp. all) players see their target in its outcome such that all strategies have a memory size of at most $n$.
An NE from $v_{0}\in V$ such that no targets are visited in its outcome in the reachability game $\mathcal{G}=(\mathcal{A},(\mathsf{Reach}(T_{i}))_{i\in\llbracket{}n\rrbracket{}})$ is an NE in the shortest-path game $(\mathcal{A},(\mathsf{SPath}^{T_{i}}_{w})_{i\in\llbracket{}n\rrbracket{}})$ for all weight functions $w$. Therefore, Theorem 5.16 can also be applied to obtain upper bounds on the memory needed to implement an NE in a shortest-path game such that that no player visits their target.
Corollary 5.19.
Let $\mathcal{G}^{\prime}=(\mathcal{A},(\mathsf{SPath}^{T_{i}}_{w})_{i\in\llbracket{}n\rrbracket{}})$ be a shortest-path game. If there exists an NE from a vertex $v_{0}$ such that no players see their target in its outcome, then there is a finite-memory NE from $v_{0}$ such that no players see their target in its outcome such that all strategies have a memory size of at most $n$.
Proof 5.20.
Let $\mathcal{G}=(\mathcal{A},(\mathsf{Reach}(T_{i}))_{i\in\llbracket{}n\rrbracket{}})$ be the reachability game obtained by ignoring weights in $\mathcal{G}^{\prime}$. Let $\sigma$ be a strategy profile and $v_{0}\in V$. Assume that $\mathsf{Out}(\sigma,v_{0})\notin\mathsf{Reach}(T_{i})$ for all $i\in\llbracket{}n\rrbracket{}$. To end the proof, it suffices to show that $\sigma$ is an NE from $v_{0}$ in $\mathcal{G}$ if and only if $\sigma$ is an NE from $v_{0}$ in $\mathcal{G}^{\prime}$ (and the result then follows from Corollary 5.18). This follows from the fact that a deviation of a player with respect to $\sigma$ is profitable in $\mathcal{G}$ if and only if it is profitable in $\mathcal{G}^{\prime}$: the cost of a player in $\mathcal{G}^{\prime}$ is finite if and only if their target is visited.
Remark 5.21.
The upcoming Theorem 5.22 provides an upper bound of $2n$ on the memory required to implement an NE in a shortest-path game with an outcome that does not visit the target of any player. By using Theorem 5.16 instead, we obtain the more precise result stated above.